# The Impact of Imperfect XAI on Human-AI Decision-Making

### v3

@@ Abstract.
Explainability techniques are rapidly being developed to improve human-AI decision-making across various cooperative work settings. Consequently, previous research has evaluated how decision-makers collaborate with imperfect AI by investigating appropriate reliance and task performance with the aim of designing more human-centered computer-supported collaborative tools. Several human-centered explainable AI (XAI) techniques have been proposed in hopes of improving decision-makers’ collaboration with AI; however, these techniques are grounded in findings from previous studies that primarily focus on the impact of incorrect AI advice. Few studies acknowledge the possibility for the explanations to be incorrect even if the AI advice is correct. Thus, it is crucial to understand how imperfect XAI affects human-AI decision-making. In this work, we contribute a robust, mixed-methods user study with 136 participants to evaluate how incorrect explanations influence humans’ decision-making behavior in a bird species identification task taking into account their level of expertise and an explanation’s level of assertiveness. Our findings reveal the influence of imperfect XAI and humans’ level of expertise on their reliance on AI and human-AI team performance. We also discuss how explanations can deceive decision-makers during human-AI collaboration. Hence, we shed light on the impacts of imperfect XAI in the field of computer-supported cooperative work and provide guidelines for designers of human-AI collaboration systems.
Human-AI Collaboration, Explainable AI, XAI for Computer Vision
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; November 09–13, 2024; San José, Costa Ricaprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Computer vision tasks
@@ 1. Introduction
Inappropriate reliance can also occur in the presence of imperfect XAI, a term that we dubbed to represent the phenomenon where an explanation can be generally correct or incorrect.  Papenmeier et al. (2019) use the term ‘explanation fidelity’ while  Kroeger et al. (2023) use the term ‘faithfulness’ to measure how “truthful” an explanation is. We use the term imperfect XAI to align with terms that already exist, such as imperfect AI, in the CSCW and human-computer interaction (HCI) communities. Specifically, we define imperfect XAI as post-hoc explanation techniques that can potentially provide explanations that are misaligned with the predicted class. We view explanation fidelity or faithfulness to be a term that can be used under the umbrella term of imperfect XAI; we view explanation fidelity to refer to the in-between cases of explanation correctness, such as when an explanation is partially correct. Imperfect XAI can exist regardless of whether the AI’s advice is correct or not. AI explanations may oversimplify or improperly estimate complex models in order to make them more interpretable, deceiving and misleading the human decision-maker. As a result, non-experts may be more prone to under- or over-relying on AI advice in the presence of incorrect explanations. Within knowledge transfer scenarios, this could cause the non-expert to learn incorrect information about a given class. Furthermore, previous research shows that the language tone within natural language explanations can impact decision-makers (Calisto et al., 2023). As it is underexplored how this affects humans’ approriate reliance when interacting with imperfect explanations, it is necessary to investigate how the communication style of explanations (e.g., the language tone, information provided, and representation) impacts human-AI collaborations across levels of expertise  (Kim and Song, 2023; Calisto et al., 2023).
@@
Collaborating with imperfect AI is not a new concept to the CSCW and HCI communities (Hemmer et al., 2021; Lindvall et al., 2021; Kocielnik et al., 2019). Despite numerous user studies over the years investigating human-AI collaborations and XAI, few have formally acknowledged the existence of imperfect XAI in their studies (Papenmeier et al., 2019; Hemmer et al., 2021). By formally acknowledging the existence of imperfect XAI in human-AI collaboration, research has several new interesting dimensions to explore. Although numerous user studies seek to understand how humans align, perceive, and interact with different types of explanations in various human-AI collaboration scenarios (e.g.,  (Chen et al., 2023; Kim et al., 2023b)), few studies explore the impact that incorrect or “noisy” explanations have on human-AI collaboration (Vasconcelos et al., 2023; Jang et al., 2023; Papenmeier et al., 2019; Hemmer et al., 2021). Recent work has used technical approaches to mitigate “noisy” or incorrect natural language explanations (Jang et al., 2023), and  Kroeger et al. (2023) propose metrics to algorithmically evaluate the effectiveness of the generated post-hoc natural language explanations. However, no studies to our knowledge investigate how the interaction between the correctness of explanations and the decision-makers’ level of expertise impact appropriate reliance on AI, human-AI team performance, and the extent to which AI explanations deceive decision-makers.
@@
  1. 1
    How does the correctness of explanations affect appropriate reliance on AI, and to what extent does the decision-maker’s level of expertise and the explanation’s assertiveness moderate this effect?
  2. 2
@@
Through a mixed-methods study, we answer our research questions by asking participants to classify bird images in two phases: without any advice from AI (phase 1) and then showing the AI’s advice and explanation (phase 2). To answer RQ 1, we design a research model based on phenomena from relevant research in CSCW and conduct rigorous moderation analyses based on (Hayes, 2017). Our analyses leverage the appropriate reliance metrics defined by  Schemmer et al. (2023). We design our study to be within-subjects for the correctness of the explanation and the assertiveness of explanations allowing us to answer RQ 2 and between-subjects for the explanation modality allowing us to answer RQ 3. Moreover, we calculate the magnitude of deception caused by incorrect explanations compared to correct explanations across both explanation modalities and levels of expertise to account for the impact of imperfect XAI on humans’ decision-making behavior. This measurement gives us insight into RQ 4. Lastly, we conduct an inductive content analysis based on Gioia et al. (2013) to assess the open-ended responses from participants to gain insight into designing for imperfect XAI in human-AI collaborations.
@@
The remainder of this article is structured as follows: First, we present related literature on imperfect AI systems, human-AI collaboration, and explainability (Section 2, p. 2) before outlining our theoretical development (Section 3, p. 3) and methodology for conducting an empirical study (Section 4, p. 4). We then present the results of our work for two different types of explanations (Section 5, p. 5). After that, we discuss the implications (Section 6, p. 6) and limitations (Section 7, p. 7) of our findings. Finally, we end our article with a brief conclusion (Section 8, p. 8).
@@ 2.1. Decision-Making with Imperfect AI/XAI
Several recent studies investigate how programmers collaborate with Copilot, an imperfect AI programming assistant (e.g.,  (Vasconcelos et al., 2023; Barke et al., 2023; Dakhel et al., 2023)). One of those studies specifically looks at how to convey the uncertainty of outputs from Copilot (Vasconcelos et al., 2023). By highlighting code that is most likely going to be edited by the programmer instead of highlighting based on the probability of the code being generated, they observe that programmers arrive at solutions faster. Furthermore,  Dakhel et al. (2023) conclude that GitHub Copilot is valuable for expert programmers, but something non-expert programmers should be cautious about.
@@ 3. Theoretical Development
Figure 1. Different paths that human decision-makers could follow based on receiving AI and XAI advice. This figure expands that presented by  Schemmer et al. (2023) by contributing the XAI advice dimension. The XAI advice is simplified into correct and incorrect explanations. The green checkmarks represent correct advice/decisions, while the red ‘x‘ represents incorrect advice/decisions.
Thus, in this work, we draw from the conceptualization on appropriate reliance previous research has established (Bansal et al., 2021; Schemmer et al., 2022, 2023). We specifically build on the conceptualization presented by  Schemmer et al. (2023) in Figure 1 by adding a new dimension to consider when investigating appropriate reliance in human-AI collaboration: The correctness of XAI advice. We simplified the correctness of XAI advice to be a binary case of correct or incorrect in Figure 1.
@@
With this new dimension for analyzing human-AI collaborations, we investigate the effect that imperfect explanations have on humans’ decision-making; we investigate this relation in a sequential decision-making scenario. Based on the constructs of relative AI reliance (RAIR) and relative self-reliance (RSR), we account for the appropriateness of reliance (Schemmer et al., 2023) [^1] Appropriateness of reliance is the quantitative measurement for appropriate reliance. These terms will be used interchangeably throughout the article.. RAIR comprises the cases in which the human corrects their initial incorrect decision by overriding it with the correct AI advice. On the other hand, RSR comprises all cases in which the human makes an initially correct decision, the AI system gives incorrect advice, and the human rightly dismisses this advice. Thus, we use appropriateness of reliance as the dependent variable in our research model (see Figure 2). In the recent work of Schoeffer et al. (2022) the authors investigate how explanations affect distributed fairness in AI-assisted decision-making. Their study shows that task-relevant explanations impact humans’ reliance behavior into increasing stereotype-based errors. We apply these findings to our research model and assume that for the cases in which the AI provides correct advice, explanations will affect RAIR. Accordingly, we hypothesize:
  1. 1
@@ 4. Methodology
In this section, we describe the task of bird species identification, the experiment design, the recruitment process of participants, and the development of the explanations we use in the study. Finally, we end this section by outlining the data we selected to use for the study and the metrics we use to analyze the results.
@@ 4.1. Task Domain: Bird Species Identification
While identifying bird species from images may not be posed as a high-stakes task in our study, this task is imperative to conserving and managing species and biodiversity (Akçay et al., 2020). Furthermore, the task of fine-grained image classification, such as bird species identification, is comparable to higher-stakes tasks, such as identifying diseases from medical imagery (Tschandl et al., 2020). For example, radiologists collaborating with an imperfect AI and imperfect XAI to diagnose diseases present in chest x-rays would go through a similar visual decision-making process as if they were trying to classify an image of a Bewick Wren in our study interface.  Kayser et al. (2022) propose an imperfect natural language explanation for chest x-rays, similar to the explanation we implement in our study, which helps bridge our findings between bird species identification and higher-stakes tasks.
Previous studies that focused on human-centered XAI and human-AI collaboration also use the domain of bird species identification to better understand human-AI collaboration (e.g.,  (Kim et al., 2023b; Nguyen et al., 2022; Cabrera et al., 2023)). However, few previous works focus on human-AI collaboration for decision-making in the wildlife conservation domain overall. Yet, the field of AI for wildlife conservation is rapidly growing (Tuia et al., 2022) and could benefit from research related to CSCW and HCI.
@@ 4.2. Study Design
To answer our research questions, we design a mixed between- and within-subjects study to examine various effects of explanations on appropriate reliance. The IRB-approved study was carefully reviewed by two experienced birders: one experienced birder is a migration counter for a bird sanctuary, and the other experienced birder holds a graduate degree in environmental science, conducted research at a nature center, and worked at a nature conservancy. The procedure of the study follows the design outlined in Figure 3 and is divided into four different parts ($A-D$), which we explain in further detail below.
@@
The study begins with part A in Figure 3: In a bird identification test, we assess participants’ expertise in classifying six different species of birds [^2] Specific birds used for the bird identification test are reported in Section A.2, p. A.2.. We distinguish the six bird images based on their level of difficulty. This level of difficulty is derived from discussions with an experienced migration counter from a bird sanctuary. While previous work has collected participants’ self-perception of expertise (Kim et al., 2023b), this method is subjective and participants may self-perceive their skills differently.  Kazemitabaar et al. (2023) measure participants “experience-level” through log data instead of subjective measures. We, similarly, try to avoid defining expertise subjectively. For the purpose of our analyses, we identify two different levels of expertise: non-experts and experts.
@@
The human-AI bird identification task (part C of Figure 3) consists of two phases. For each treatment condition, the participants are asked to initially identify the bird species from an image (phase 1 in Figure 4). After submitting an initial identification, they are shown the AI’s prediction along with the explanation, and again have to submit an identification for the bird species in the image (phase 2 in Figure 4). The structure of phases one and two are corroborated with previous work (Green and Chen, 2019). Initially, participants must click on a button that shows “Show AI Explanation”. Without revealing the explanation, the participant cannot proceed to the next question. This is one way for us to ensure that the participant acknowledges the presence of an explanation. Overall, participants do this process for twelve different random bird images.
@@
After finishing the task, participants must fill out an additional questionnaire (part D of Figure 3). Here, we qualitatively assess participants’ ability to properly rely on the AI based on the explanations that they were shown. Thus we ask them: “Under what circumstances would you prefer assertive (e.g., “definitely”, “clearly”) versus non-assertive (e.g., “might be”, “appears to be”) versus neutral explanations and why?”. Aside from this open-text question, we also ask participants about their occupation and the regions in North America that they are most familiar with in terms of bird species.
@@ 4.3. Data Selection
We create a dataset of bird images and explanations to show participants by manually curating bird images from the well-established CUB-200-2011 (Wah et al., 2011) dataset and explanations from the Generating Visual Explanations model (Hendricks et al., 2021). The original dataset consists of 11,788 images of 200 different bird species and is split into 5,994 training and 5,794 test images. Each bird species is represented with around 60 images of the respective bird class. When curating birds, we first filtered for bird families with several species in the CUB dataset. We specifically filtered out every bird class that is not a part of the Warblers, Wrens, Swallows, Sparrows, or Finches/Grosbeaks families. After applying this filter, we had $1,864$ out of $5,794$ images from the test set of CUB-200-2011. Of those $1,864$ images, $1609$ were predicted correctly by the AI, and $255$ images were predicted incorrectly by the model.
After filtering the bird species, multiple researchers on our team separately classified the natural language explanations and the visual, example-based explanations for a subset of the $1,864$ birds as incorrect or correct. Cases of doubt were discussed by a subset of the research team and excluded from consideration if an agreement was not met. In total, we identified 10 examples for each category [^4] The four categories are identified in Section 4.2. and explanation type. In some cases, the example-based and natural language explanations for a single bird are used. As a result, the dataset represents 66 different images and 43 different bird species from the CUB-200-2011 dataset.
@@
Figure 5. Representative examples of the example-based and natural language explanations for each scenario: CC, CI, IC, and II. The class of the example-based images in the explanation is not shown to participants during the study. The red and green coloring on the natural language explanations was also not shown during the study. This is only provided in the figure to guide the reader. The natural language explanation for the Cerulean Warbler is incorrect because this bird species does not have a grey head. The natural language explanation for the Yellow-billed Cuckoo is incorrect because this bird species is brown with a white belly, has a gold and black beak, and does not have a black wing.
For the example-based explanations, we define a correct explanation as the three most similar images belonging to the predicted class (as shown in phase 2 of Figure 4). We define an incorrect explanation to be most similar to at least one image that is not of the predicted class. This means that incorrect example-based explanations incorporate logical errors as the examples shown are dissimilar from the predicted class. Moreover, such incorrect explanations can inhold an inconsistency as the shown examples might differ in the classes shown. However, we only choose to show participants incorrect explanations that have at least two images that are not of the predicted class. For example, in Figure 5, the AI correctly predicts a Nashville Warbler; however, the three most similar examples are a Painted Bunting, a Yellow-Breasted Chat, and an American Redstart. In some cases where the advice is correct and the explanation is incorrect [^5] The explanation does not align with the predicted class., the explanation may align with the ground truth class. For example, for a Tennessee Warbler, the AI predicts Orange-crowned Warbler (incorrect advice since the false bird species is predicted), but the three most similar examples are all of Tennessee Warblers (incorrect explanation since the examples’ bird species do not align with prediction). It’s possible that a model could be relying on spurious patterns to make classifications (Plumb et al., 2021). Since we are dealing with an imperfect AI, we do not choose to exclude such cases from our dataset.

### v4

@@ Abstract.
Explainability techniques are rapidly being developed to improve human-AI decision-making across various cooperative work settings. Consequently, previous research has evaluated how decision-makers collaborate with imperfect AI by investigating appropriate reliance and task performance with the aim of designing more human-centered computer-supported collaborative tools. Several human-centered explainable AI (XAI) techniques have been proposed in hopes of improving decision-makers’ collaboration with AI; however, these techniques are grounded in findings from previous studies that primarily focus on the impact of incorrect AI advice. Few studies acknowledge the possibility of the explanations being incorrect even if the AI advice is correct. Thus, it is crucial to understand how imperfect XAI affects human-AI decision-making. In this work, we contribute a robust, mixed-methods user study with 136 participants to evaluate how incorrect explanations influence humans’ decision-making behavior in a bird species identification task, taking into account their level of expertise and an explanation’s level of assertiveness. Our findings reveal the influence of imperfect XAI and humans’ level of expertise on their reliance on AI and human-AI team performance. We also discuss how explanations can deceive decision-makers during human-AI collaboration. Hence, we shed light on the impacts of imperfect XAI in the field of computer-supported cooperative work and provide guidelines for designers of human-AI collaboration systems.
Human-AI Collaboration, Explainable AI, Explainable AI for Computer Vision
copyright: rightsretainedjournal: PACMHCIjournalyear: 2024journalvolume: 8journalnumber: CSCW1article: 183publicationmonth: 4doi: 10.1145/3641022ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Computer vision tasks
@@ 1. Introduction
Inappropriate reliance can also occur in the presence of imperfect XAI, a term that we introduce to represent the phenomenon where an explanation reveals evidence that does not necessarily comply with the prediction.  Papenmeier et al. (2019) use the term ‘explanation fidelity’ while  Kroeger et al. (2023) use the term ‘faithfulness’ to measure how “truthful” an explanation is. We use imperfect XAI to align with existing terms, such as imperfect AI, in the CSCW and human-computer interaction (HCI) communities. Specifically, we define imperfect XAI as explanation techniques that can potentially provide explanations that do not fit with the AI’s predictions. We view explanation fidelity or faithfulness as a term that can be used under the umbrella term of imperfect XAI; we view explanation fidelity as referring to the continuum of explanation correctness, such as when explanations are partially correct. Imperfect XAI can exist regardless of whether the AI’s advice is correct or not. AI explanations may oversimplify or improperly estimate complex models in order to make them more interpretable, deceiving and misleading the human decision-maker. As a result, non-experts may be more prone to under- or over-relying on AI advice in the presence of incorrect explanations. Within knowledge transfer scenarios, this could cause the non-expert to learn incorrect information about a given class. Furthermore, previous research shows that the language tone within natural language explanations can impact decision-makers (Calisto et al., 2023). However, the effect of language tone on humans’ appropriate reliance when interacting with imperfect explanations is underexplored. Therefore, it is necessary to investigate how the communication style of explanations (e.g., the language tone, the information provided, and their representations) impacts human-AI collaborations across levels of expertise  (Kim and Song, 2023; Calisto et al., 2023).
@@
Collaborating with imperfect AI is not a new concept to the CSCW and HCI communities (Hemmer et al., 2021; Lindvall et al., 2021; Kocielnik et al., 2019). Despite numerous user studies over the years investigating human-AI collaborations and XAI, few have formally acknowledged the existence of imperfect XAI in their studies (Papenmeier et al., 2019; Hemmer et al., 2021). By formally acknowledging the existence of imperfect XAI in human-AI collaboration, research has several new interesting dimensions to explore. Although numerous user studies seek to understand how humans align, perceive, and interact with different types of explanations in various human-AI collaboration scenarios (e.g.,  (Chen et al., 2023; Kim et al., 2023b)), few studies explore the impact that incorrect or “noisy” explanations have on human-AI collaboration (Vasconcelos et al., 2023; Jang et al., 2023; Papenmeier et al., 2019; Hemmer et al., 2021). Recent work has used technical approaches to mitigate “noisy” or incorrect natural language explanations (Jang et al., 2023), and  Kroeger et al. (2023) propose metrics to algorithmically evaluate the effectiveness of the generated post-hoc natural language explanations. However, to our knowledge, no studies investigate how the interaction between the correctness of explanations and the decision-maker’s level of expertise impact appropriate reliance on AI, human-AI team performance, and the extent to which AI explanations deceive decision-makers.
@@
  1. RQ1:
    How does the correctness of explanations affect appropriate reliance on AI, and to what extent do the decision-maker’s level of expertise and the explanation’s assertiveness moderate this effect?
  2. RQ2:
@@ 4. RQ4:
Through a mixed-methods study, we answer our research questions by asking participants to classify bird images in two phases: without any advice from AI (phase 1) and then showing the AI’s advice and explanation (phase 2). To answer RQ 1, we design a research model based on phenomena from relevant research in CSCW and conduct rigorous moderation analyses based on (Hayes, 2017). Our analyses leverage the appropriate reliance metrics defined by  Schemmer et al. (2023b). We design our study to be within-subjects for the correctness of the explanation and the assertiveness of explanations allowing us to answer RQ 2 and between-subjects for the explanation modality allowing us to answer RQ 3. Moreover, we calculate the magnitude of deception caused by incorrect explanations compared to correct explanations across both explanation modalities and levels of expertise to account for the impact of imperfect XAI on humans’ decision-making behavior. This measurement gives us insight into RQ 4. Lastly, we conduct an inductive content analysis based on Gioia et al. (2013) to assess the open-ended responses from participants to gain insight into designing for imperfect XAI in human-AI collaborations.
@@
The remainder of this article is structured as follows: We present related literature on imperfect AI systems, human-AI collaboration, and explainability (Section 2, p. 2) before outlining our theoretical development (Section 3, p. 3) and methodology for conducting an empirical study (Section 4, p. 4). We then present the results of our work for two different types of explanations (Section 5, p. 5). After that, we discuss the implications (Section 6, p. 6) and limitations (Section 7, p. 7) of our findings. Finally, we end our article with a brief conclusion (Section 8, p. 8).
@@ 2.1. Decision-Making with Imperfect AI/XAI
Several recent studies investigate how programmers collaborate with Copilot, an imperfect AI programming assistant (e.g.,  (Vasconcelos et al., 2023; Barke et al., 2023; Dakhel et al., 2023)). One of those studies specifically looks at how to convey the uncertainty of outputs from Copilot (Vasconcelos et al., 2023). By highlighting code that is most likely going to be edited by the programmer instead of highlighting based on the probability of the code being generated, they observe that programmers arrive at solutions faster. Furthermore,  Dakhel et al. (2023) conclude that GitHub Copilot is valuable for expert programmers but something non-expert programmers should be cautious about.
@@ 3. Theoretical Development
Figure 1. Different paths that human decision-makers could follow based on receiving AI and XAI advice. This figure expands that presented by  Schemmer et al. (2023b) by contributing the XAI advice dimension. The XAI advice is simplified into correct and incorrect explanations. The green checkmarks represent correct advice/decisions, while the red ‘x‘ represents incorrect advice/decisions.
Thus, in this work, we draw from the conceptualization on appropriate reliance previous research has established (Bansal et al., 2021; Schemmer et al., 2022, 2023b; Schoeffer et al., 2023). We specifically build on the conceptualization presented by  Schemmer et al. (2023b) in Figure 1 by adding a new dimension to consider when investigating appropriate reliance in human-AI collaboration: The correctness of XAI advice. We simplified the correctness of XAI advice to be a binary case of correct or incorrect in Figure 1.
@@
With this new dimension for analyzing human-AI collaborations, we investigate the effect that imperfect explanations have on humans’ decision-making; we investigate this relation in a sequential decision-making scenario. Based on the constructs of relative AI reliance (RAIR) and relative self-reliance (RSR), we account for the appropriateness of reliance (Schemmer et al., 2023b) [^1] Appropriateness of reliance is the quantitative measurement for appropriate reliance. These terms will be used interchangeably throughout the article.. RAIR comprises the cases in which the human corrects their initially incorrect decision by overriding it with the correct AI advice. On the other hand, RSR comprises all cases in which the human initially makes a correct decision, the AI system gives incorrect advice, and the human rightly dismisses this advice. Thus, we use appropriateness of reliance as the dependent variable in our research model (see Figure 2). In the recent work of Schoeffer et al. (2024) the authors investigate how explanations affect distributed fairness in AI-assisted decision-making. Their study shows that task-relevant explanations impact humans’ reliance behavior into increasing stereotype-based errors. We apply these findings to our research model and assume that for the cases in which the AI provides correct advice, explanations will affect RAIR. Accordingly, we hypothesize:
  1. Hypothesis 1:
@@ 4. Methodology
In this section, we describe the task of bird species identification, the experiment design, the recruitment process of participants, and the development of the explanations we use in the study. Finally, we end this section by outlining the data we select to use for the study and the metrics we use to analyze the results.
@@ 4.1. Task Domain: Bird Species Identification
While identifying bird species from images may not be posed as a high-stakes task in our study, this task is imperative to conserving and managing species and biodiversity (Akçay et al., 2020). Furthermore, the task of fine-grained image classification, such as bird species identification, is comparable to higher-stakes tasks, such as identifying diseases from medical imagery (Tschandl et al., 2020). For example, radiologists collaborating with an imperfect AI and imperfect XAI to diagnose diseases present in chest X-rays would go through a similar visual decision-making process as if they were trying to classify an image of a Bewick Wren in our study interface.  Kayser et al. (2022); Hou et al. (2021) propose an imperfect natural language explanation for chest x-rays, similar to the explanation we implement in our study, which helps bridge our findings between bird species identification and higher-stakes tasks.
Previous studies that focused on human-centered XAI and human-AI collaboration also use the domain of bird species identification to understand better human-AI collaboration (e.g.,  (Kim et al., 2023b; Nguyen et al., 2022; Cabrera et al., 2023)). However, few previous works focus on human-AI collaboration for decision-making in the wildlife conservation domain overall. Yet, the field of AI for wildlife conservation is rapidly growing (Tuia et al., 2022) and could benefit from research related to CSCW and HCI.
@@ 4.2. Study Design
To answer our research questions, we design a mixed between- and within-subjects study to examine various effects of explanations on appropriate reliance. Our study was carefully reviewed by two experienced birders: one experienced birder is a migration counter for a bird sanctuary, and the other experienced birder holds a graduate degree in environmental science, conducted research at a nature center, and worked at a nature conservancy. The procedure of the study follows the design outlined in Figure 3 and is divided into four different parts ($A-D$), which we explain in further detail below.
@@
The study begins with part A in Figure 3: In a bird identification test, we assess participants’ expertise in classifying six different species of birds [^2] Specific birds used for the bird identification test are reported in Section A.2, p. A.2.. We distinguish the six bird images based on their level of difficulty. This level of difficulty is derived from discussions with an experienced migration counter from a bird sanctuary. While previous work has collected participants’ self-perception of expertise (Kim et al., 2023b), this method is subjective, and participants may self-perceive their skills differently.  Kazemitabaar et al. (2023) measure participants “experience-level” through log data instead of subjective measures. Similarly, we try to avoid defining expertise subjectively. For the purpose of our analyses, we identify two different levels of expertise: non-experts and experts.
@@
The human-AI bird identification task (part C of Figure 3) consists of two phases. For each treatment condition, the participants are asked to initially identify the bird species from an image (phase 1 in Figure 4). After submitting an initial identification, they are shown the AI’s prediction along with the explanation, and again, they have to submit an identification for the bird species in the image (phase 2 in Figure 4). The structure of phases one and two are corroborated with previous work (Green and Chen, 2019). Initially, participants must click on a button that shows “Show AI Explanation”. Without revealing the explanation, the participant cannot proceed to the next question. This is one way for us to ensure that the participant acknowledges the presence of an explanation. Overall, participants do this process for twelve different random bird images.
@@
After finishing the task, participants must fill out an additional questionnaire (part D of Figure 3). Here, we qualitatively assess participants’ ability to properly rely on the AI based on the explanations that they were shown. Thus we ask them: “Under what circumstances would you prefer assertive (e.g., “definitely”, “clearly”) versus non-assertive (e.g., “might be”, “appears to be”) versus neutral explanations and why?”. Aside from this open-text question, we also ask participants about their occupations and the regions in North America that they are most familiar with in terms of bird species.
@@ 4.3. Data Selection
We create a dataset of bird images and explanations to show participants by manually curating bird images from the well-established CUB-200-2011 (Wah et al., 2011) dataset and explanations from the Generating Visual Explanations model (Hendricks et al., 2021). The original dataset consists of 11,788 images of 200 different bird species and is split into 5,994 training and 5,794 test images. Each bird species is represented with around 60 images of the respective bird class. When curating birds, we first filtered for bird families with several species in the CUB dataset. We specifically filtered out every bird class that is not a part of the Warblers, Wrens, Swallows, Sparrows, or Finches/Grosbeaks families. After applying this filter, we had $1,864$ out of $5,794$ images from the test set of CUB-200-2011. Of those $1,864$ images, $1,609$ were predicted correctly by the AI, and $255$ images were predicted incorrectly by the model.
After filtering the bird species, multiple researchers on our team separately classified the natural language explanations and the visual, example-based explanations for a subset of the $1,864$ birds as incorrect or correct. Cases of doubt were discussed by a subset of the research team and excluded from consideration if an agreement was not met. In total, we identified ten examples for each category [^4] The four categories are identified in Section 4.2. and explanation type. In some cases, the example-based and natural language explanations for a single bird are used. As a result, the dataset represents 66 different images and 43 different bird species from the CUB-200-2011 dataset.
@@
Figure 5. Representative examples of the example-based and natural language explanations for each scenario: CC, CI, IC, and II. The class of the example-based images in the explanation is not shown to participants during the study. The red and green coloring on the natural language explanations was not shown during the study. This is only provided in the figure to guide the reader. The natural language explanation for the Cerulean Warbler is incorrect because this bird species does not have a grey head. The natural language explanation for the Yellow-billed Cuckoo is incorrect because this bird species is brown with a white belly, has a gold and black beak, and does not have a black wing.
For the example-based explanations, we define a correct explanation as the three most similar images belonging to the predicted class (as shown in phase 2 of Figure 4). We define an incorrect explanation to be most similar to at least one image that is not of the predicted class. This means that incorrect example-based explanations incorporate logical errors as the examples shown are dissimilar from the predicted class. Moreover, such incorrect explanations can have an inconsistency as the examples shown might differ in the classes shown. However, we only choose to show participants incorrect explanations that have at least two images that are not of the predicted class. For example, in Figure 5, the AI correctly predicts a Nashville Warbler; however, the three most similar examples are a Painted Bunting, a Yellow-Breasted Chat, and an American Redstart. In some cases where the advice is correct and the explanation is incorrect [^5] The explanation does not align with the predicted class., the explanation may align with the ground truth class. For example, for a Tennessee Warbler, the AI predicts an Orange-crowned Warbler (incorrect advice since the wrong bird species is predicted), but the three most similar examples are all of Tennessee Warblers (incorrect explanation since the examples’ bird species do not align with prediction). It’s possible that a model could be relying on spurious patterns to make classifications (Plumb et al., 2021). Since we are dealing with an imperfect AI, we do not choose to exclude such cases from our dataset.